# A unified framework for geometry-independent operator learning in cardiac electrophysiology simulations

### v2

@@ 3.3 Toward domain-invariant generalisation
Prior studies have evaluated model performance primarily under variations in pacing site [yin2024scalable]. In contrast, a central contribution of our work is demonstrating that a single model can generalise across both diverse pacing locations and simulation parameters. This capability is essential for clinical applicability, since the pacing site is often variable or unknown in practice. Our results also show that the unified model, trained jointly across all seven pacing sites, consistently and significantly outperforms models trained for individual pacing locations. The unified model achieves a 16.5% reduction in mean MAE, and a 1.6% increase in average SSIM. This demonstrates that our model effectively extracts generalizable electrophysiological principles that are independent of the specific pacing location, making it a robust and scalable solution for real-world clinical applications where prior knowledge of the stimulation site is unavailable.
Our cross-domain experiments reveal a critical insight. The apparent limitation of single-domain models is not a shortcoming of the ViT architecture, but a reflection of latent heterogeneity in clinical imaging data. By explicitly combining datasets from distinct acquisition and annotation protocols, we demonstrate that the model naturally discards protocol-specific shortcuts and focuses on the fundamental anatomical structure. This approach not only rescues performance on external cohorts but also slightly improves internal predictions.
This finding establishes a generalizable strategy for achieving domain-invariant medical image analysis. The universal coordinate projection provides the necessary geometry-invariant foundation, while training on heterogeneous data acts as a regularisation method in feature space, compelling the model to learn only the biophysically relevant structures common to both domains and emphasising that true translational power derives from integrating diverse, multi-centre data rather than from isolated single-site optimisation.
@@ 3.4 Limitations and future directions
Beyond feature representation, limitations also arise from the use of intrinsic anatomical coordinate systems. For the left atrium, the UAC system maps a closed, topologically complex three-dimensional surface onto a flat two-dimensional domain. This flattening necessarily requires introducing a cut line on the surface, which appears as a seam in the coordinate representation. As a result, points that are physically adjacent on the original anatomy become separated across this seam, creating an artificial spatial discontinuity for the learning model. Consistent with this construction, our results reveal a narrow, coherent band of elevated prediction error aligned with the UAC seam, indicating that the neural operator struggles to interpolate smoothly across this boundary.
@@
In addition to seam-related effects, further localised errors are observed in regions of pronounced anatomical and electrophysiological complexity, particularly around the pulmonary vein junctions and the left atrial appendage. These regions exhibit substantial inter-patient variability that cannot be fully normalised by the UAC projection, leading to residual geometric heterogeneity in the coordinate domain. Moreover, activation wavefronts in these areas often display steep local gradients in activation time, increasing sensitivity to small spatial misalignments. Together, these factors likely contribute to the small but systematic clusters of elevated error observed in physiologically critical regions. These observations indicate that while the proposed framework captures global activation patterns and long-range structure effectively, resolving fine-scale, anatomy-driven variations remains challenging.
@@
By contrast, ventricular geometries parameterised using the UVC do not exhibit a comparable global seam-induced discontinuity. Instead, elevated errors are primarily confined to regions introduced by geometric truncation, such as the basal valve plane and the interface with the right ventricle. This contrast highlights that coordinate-induced limitations are strongly dependent on underlying anatomical topology rather than intrinsic to the operator-learning framework itself, and that intrinsic coordinate representations may behave differently across cardiac chambers.
@@
Finally, the framework was trained exclusively on high-fidelity simulated datasets. Although these simulations span a wide range of anatomies and electrophysiological parameters, validation against clinical electroanatomic mapping data will be essential to assess robustness under real-world conditions. In addition, the current implementation focuses primarily on atrial and ventricular geometries in isolation; extending the approach to whole-heart or fully biventricular modelling will require adapting the anatomical parameterisation and scaling the data generation process. Future directions include integrating hybrid simulation–clinical datasets, refining anatomical representations in regions of high complexity, and exploring adaptive spatial resolution strategies to further improve local accuracy without compromising inference efficiency.
@@ 4 Methods
We describe a computational framework that combines high-fidelity electrophysiology simulations with neural operator learning to enable efficient prediction of LAT maps (Fig. 4).
Figure 4: The overview of our proposed computational framework for operator learning for electrophysiology (EP) simulations. a, Preparation of left atrium (LA) mesh, mapping the 3D vertex coordinates defining the mesh to a 2D grid by Universal Atrial Coordinates (UAC), and identifying the pacing locations. b, The application of our GPU-based FEM solver to produce a large amount of simulated local activation time (LAT) maps, which are also projected to 2D grids by UAC. c, The training of neural operators using the structured simulation data.
4.1 Image data processing
4.1.1 Datasets description
@@
Notably, Cohorts A and B were derived from two distinct and independent studies conducted by separate research groups operating within independent healthcare systems. This independence in data acquisition and analysis supports robust cross-domain evaluation. Moreover, despite adhering to equivalent segmentation and processing workflows [solis2023evaluation, razeghi2020cemrgapp], the two cohorts exhibit meaningful domain shifts. Variations arise from MRI acquisition parameters (scanner vendor, pulse sequences, and field strength), from different practices in anatomical trimming (particularly of pulmonary veins and the LA appendage), and from inter-operator differences in anatomical landmarking and boundary definitions. These protocol-level and anatomical choices shape the resulting geometries and fibrosis patterns, such that models trained exclusively on Cohort A typically encounter Cohort B as an out-of-distribution domain. These differences are not physiological changes. These differences are not physiological. They are artefacts of data preparation, yet they change the geometric domain on which the operator is defined.
4.1.2 Universal coordinates projection
Spatial coordinates, fibre directions, and LAT fields from patient-specific mesh geometries are projected onto structured grids of fixed resolution using universal anatomical coordinate systems. For the left atrium, UAC defines a bijective transformation
@@ 4.1.2 Universal coordinates projection
4.1.3 Identify pacing locations
@@
Pacing locations are defined across all patient anatomies using the UAC system. For each anatomical region, a 2D bounding box within the UAC domain was specified, delineating a subset of mesh vertices. The physical pacing centre is computed as the mean position of all vertices within that UAC-defined box. The stimulation region is subsequently defined by all mesh vertices located within a fixed Euclidean radius of $2.0\,\text{mm}$ from this central point, ensuring comparable stimulation volumes across all seven sites. The spatial distribution of these seven sites on a representative LA mesh is mapped onto the UAC coordinate system in Fig. 5.
@@
4.2 GPU-based FEM EP simulation
To generate a large-scale, comprehensive dataset for training, we simulated LA electrical propagation using the widely adopted monodomain equation [geselowitz1983bidomain] (further implementation details are provided in Supplementary Section 1, with an example of simulated wavefront propagation shown in Supplementary Fig. 1 and benchmark verification results in Supplementary Fig. 2). This biophysical model captures the electrical conduction through anisotropic cardiac tissue, balancing capacitive and ionic currents with diffusive current flow, governed by the longitudinal ($\sigma_{l}$) and transverse ($\sigma_{t}$) conductivities.
@@ 4.2 GPU-based FEM EP simulation
4.3 Neural operator learning
We formulate LAT map prediction as a neural operator learning problem, in which the goal is to approximate a mapping from patient-specific anatomical and electrophysiological parameters to the corresponding spatial distribution of activation times. Formally, we aim to learn an operator
@@ 4.3 Neural operator learning
where $A$ denotes the atrial geometry, $\sigma$ the conductivity tensor, $p$ the pacing site location, and $\mathrm{LAT}(\mathbf{x})$ the activation time at position $\mathbf{x}\in\Omega$ where $\Omega$ is the computational c. This framework requires the generation of large and diverse training datasets that pair input configurations $(A,\sigma,p)$ with ground-truth LAT maps obtained from high-fidelity EP simulations.
4.3.1 Model architecture
@@
Multi-modal feature integration.
The input to the neural operator is a multi-modal, multi-channel grid that encodes patient-specific mesh properties and simulation parameters into a fixed-size ten-channel representation (Fig. 6a), which can be divided into two groups. The first group comprises the seven geometric and anatomical channels, including the UAC-projected 3D vertex coordinates $(x,y,z)$ and the 3D fibre direction vectors (six channels in total, each forming a $(3,50,50)$ feature map). This is supplemented by the total surface area of the left atrium, a global anatomical feature that is broadcast across the spatial domain to form a $(1,50,50)$ constant grid. The second group comprises the two parametric features including the two conduction velocities ($\sigma_{l}$ and $\sigma_{t}$), which are also broadcast across the spatial domain to form a $(2,50,50)$ feature map, and the encoded pacing location in a $(1,50,50)$ feature map. The output of the neural network is trained against the target LAT map of shape $(1,50,50)$.
ViT-Based Spatial Representation Learning.
To model spatial dependencies across the input domain, we employ a Vision Transformer architecture that encodes the multi-channel input grid into a sequence of contextualised latent representations. A convolutional patch embedding layer [wu2021cvt] first divides the input grid into non-overlapping patches and projects each patch into a shared embedding space, producing a sequence of latent tokens. Learnable positional embeddings are added to retain spatial structure within the tokenised representation. This sequence is then processed by a stack of standard ViT blocks, each comprising multi-head self-attention and feed-forward sublayers, which enables the integration of global context and the capture of long-range spatial interactions (Fig. 6c).
Following the ViT-based encoder, the latent tokens are reshaped into a compact 2D feature map. To recover full spatial resolution, we employ a lightweight decoder consisting of a transposed convolution layer for upsampling, followed by a convolutional projection head that produces the final output map. The transposed convolution operation, originally introduced in deep learning by [zeiler2011adaptive] for adaptive deconvolutional networks and later popularised for CNN visualisation by [zeiler2014visualizing], enables learnable upsampling by applying the transpose of the convolution operator. This decoding step ensures that the global features encoded by the transformer are effectively translated back into a high-resolution spatial field for pixel-wise prediction.
4.3.2 Sobolev Loss with Spatial Smoothness Regularisations

### v3

@@ 3.3 Toward domain-invariant generalisation
Generalisation across heterogeneous input conditions is a central challenge for data-driven modelling on complex anatomical domains. Prior studies have primarily evaluated performance under variations in pacing site [yin2024scalable]. In contrast, our results show that a single model can generalise across both diverse pacing locations and simulation parameters. Models trained jointly across all pacing sites consistently outperform those trained for individual locations, indicating that the learned representations capture electrophysiological structure that is not tied to a specific stimulation configuration.
Beyond pacing variability, our cross-domain experiments reveal a more fundamental limitation. Performance degradation when models are trained on a single dataset and evaluated on an independent cohort is not primarily attributable to the choice of network architecture, but rather reflects latent heterogeneity in imaging, annotation, and preprocessing protocols. By explicitly combining datasets acquired under distinct conditions, the model is encouraged to discard protocol-specific shortcuts and instead focus on anatomical features that are consistent across domains. This strategy not only improves performance on external cohorts but also yields modest gains on internal test sets.
Together, these observations suggest a general strategy for achieving domain-invariant learning in medical imaging and simulation-based modelling. Universal coordinate representations provide a geometry-invariant foundation, while training on heterogeneous data acts as an implicit regularisation in feature space, promoting representations that reflect underlying biophysical structure rather than dataset-specific artefacts. More broadly, these results indicate that robust translational performance is more likely to arise from integrating diverse, multi centre data than from optimising models within a single acquisition domain.
@@ 3.4 Limitations and future directions
One limitation revealed by our ablation studies concerns the role of fibre orientation features. Removing fibre orientation may initially appear counterintuitive, yet it consistently improved predictive performance. Two factors are likely to contribute to this outcome. First, the current fibre orientation representation may introduce noise or uncertainty that interferes with learning the underlying electrophysiological dynamics. Second, the model appears to rely more effectively on the spatial and geometric priors provided by the anatomical coordinates and surface area features, which may already encode sufficient structural information. In this setting, explicit fibre orientation inputs may therefore be redundant or even detrimental in their present form.
@@
Beyond feature representation, limitations also arise from the anatomical coordinate parameterisation itself. For the left atrium, the UAC system maps a closed, topologically complex three-dimensional surface onto a flat two-dimensional domain, which necessarily introduces a cut line that appears as a seam in the coordinate representation. As a result, points that are physically adjacent on the original anatomy become separated across this seam, creating an artificial spatial discontinuity for the learning model. Consistent with this construction, our results reveal a narrow band of elevated prediction error aligned with the UAC seam, indicating difficulty in interpolating smoothly across this boundary.
@@
Additional localised errors are observed in regions of pronounced anatomical and electrophysiological complexity, particularly around the pulmonary vein junctions and the left atrial appendage. These regions exhibit interpatient variability that cannot be fully normalised by the UAC projection, leading to residual geometric heterogeneity in the coordinate domain. Together, these factors contribute to small but systematic clusters of elevated error in physiologically critical regions. While the proposed framework captures global activation patterns and structure over long spatial scales, resolving fine-scale anatomical variation remains challenging.
@@
By contrast, ventricular geometries parameterised using the UVC do not exhibit a comparable global seam-induced discontinuity. Instead, ventricular error behaviour is governed primarily by boundary effects arising from geometric truncation rather than by coordinate-induced artefacts. Importantly, these truncations preserve the overall ventricular topology and continuity of the domain. This contrast highlights that limitations associated with intrinsic coordinate representations depend on underlying anatomical topology rather than being intrinsic to the operator learning framework itself, and that coordinate-based representations may therefore behave differently across cardiac chambers.
@@
Finally, the framework was trained exclusively on high-fidelity simulated datasets. Although these simulations span a wide range of anatomies and electrophysiological parameters, validation against clinical electroanatomic mapping data will be essential to assess robustness under real-world conditions. In addition, the current implementation focuses primarily on atrial and ventricular geometries in isolation. Extending the approach to whole-heart or fully biventricular modelling will require adapting the anatomical parameterisation and scaling the data generation process. Future directions include integrating hybrid simulation–clinical datasets, refining anatomical representations in regions of high complexity, and exploring adaptive spatial resolution strategies to improve local accuracy while preserving computational efficiency and physical fidelity.
@@ 4 Methods
4.1 Geometry-independent operator learning framework
The proposed framework separates operator learning from geometric representation by introducing an intrinsic coordinate domain that serves as the sole space in which learning is performed. Both input fields and target outputs are represented on this intrinsic domain at a fixed resolution, which allows standard neural operator architectures defined on regular grids to be used directly. The original physical mesh is not used during learning itself and is used only to project quantities that depend on geometry and physical parameters into the intrinsic domain.
The framework consists of three stages. First, a problem-specific intrinsic coordinate map is used to project geometric coordinates, boundary information, and physical parameters from a mesh domain to a structured coordinate grid. Second, high-fidelity simulations are performed on the original mesh, and the resulting output fields are projected to the same coordinate grid to form paired training data, which are then used to train a neural operator in the coordinate domain. Third, predictions produced on the coordinate grid are interpolated back to the original mesh to enable anatomically faithful visualisation and downstream analysis.
In this study, cardiac electrophysiology serves as the primary use case, with the target field being the local activation time map (Fig. 4). An additional kinematic deformation problem is used to demonstrate that the proposed framework can be applied unchanged to mechanically governed systems with different physics and outputs.
Figure 4: Overview of the geometry-independent operator learning framework, shown for cardiac electrophysiology. a, Preparation of left atrium (LA) mesh, mapping the 3D vertex coordinates defining the mesh to a 2D grid by Universal Atrial Coordinates (UAC), and identifying the pacing locations. b, High-fidelity simulation on the original mesh and projection of output fields to the same coordinate grid to form training pairs. c, Neural operator learning in the coordinate domain and reconstruction of predictions on the original mesh for evaluation and downstream analysis.
4.2 Instantiation for cardiac electrophysiology
4.2.1 Geometries and cohorts
@@
Notably, Cohorts A and B were derived from two distinct and independent studies conducted by separate research groups operating within independent healthcare systems. This independence in data acquisition and analysis supports robust cross-domain evaluation. Moreover, despite adhering to equivalent segmentation and processing workflows [solis2023evaluation, razeghi2020cemrgapp], the two cohorts exhibit meaningful domain shifts. Variations arise from MRI acquisition parameters (scanner vendor, pulse sequences, and field strength), from different practices in anatomical trimming (particularly of pulmonary veins and the LA appendage), and from inter-operator differences in anatomical landmarking and boundary definitions. These protocol-level and anatomical choices shape the resulting geometries and fibrosis patterns, such that models trained exclusively on Cohort A typically encounter Cohort B as an out-of-distribution domain. These differences are not physiological changes. They are artefacts of data preparation, yet they change the geometric domain on which the operator is defined.
4.2.2 Intrinsic coordinate projection and reconstruction
The framework assumes the availability of an intrinsic coordinate mapping that provides a consistent parameterisation across heterogeneous geometries. Using such a mapping, spatial coordinates, fibre directions, and scalar fields defined on patient-specific meshes are projected onto structured grids of fixed resolution. For the left atrium, UAC defines a bijective transformation
@@ 4.2.2 Intrinsic coordinate projection and reconstruction
4.2.3 High-fidelity electrophysiology simulations
To generate a large-scale, comprehensive dataset for training, we simulated LA electrical propagation using the widely adopted monodomain equation [geselowitz1983bidomain] (more implementation details are provided in Supplementary Section 1, with an example of simulated wavefront propagation shown in Supplementary Fig. 1 and benchmark verification results in Supplementary Fig. 2). This biophysical model captures the electrical conduction through anisotropic cardiac tissue, balancing capacitive and ionic currents with diffusive current flow, governed by the longitudinal ($\sigma_{l}$) and transverse ($\sigma_{t}$) conductivities.
@@ 4.2.3 High-fidelity electrophysiology simulations
Pacing locations are defined across all patient anatomies using the UAC system. For each anatomical region, a 2D bounding box within the UAC domain was specified, delineating a subset of mesh vertices. The physical pacing centre is computed as the mean position of all vertices within that UAC-defined box. The stimulation region is subsequently defined by all mesh vertices located within a fixed Euclidean radius of $2.0\,\text{mm}$ from this central point, ensuring comparable stimulation volumes across all seven sites. The spatial distribution of these seven sites on a representative LA mesh is mapped onto the UAC coordinate system (Fig. 5).
@@
4.2.4 Operator learning
Within the proposed framework, we formulate LAT map prediction as a neural operator learning problem, in which the goal is to approximate a mapping from patient-specific anatomical and electrophysiological parameters to the corresponding spatial distribution of activation times. Formally, we aim to learn an operator
@@ 4.2.4 Operator learning
where $A$ denotes the atrial geometry, $\sigma$ the conductivity tensor, $p$ the pacing site location, and $\mathrm{LAT}(\mathbf{x})$ the activation time at position $\mathbf{x}\in\Omega$ where $\Omega$ is the computational domain. This framework requires the generation of large and diverse training datasets that pair input configurations $(A,\sigma,p)$ with ground-truth LAT maps obtained from high-fidelity EP simulations.
@@
Multimodal feature integration.
The neural operator operates on a grid representation with multiple channels, which jointly encode mesh geometry and simulation parameters (Fig. 6a). These channels can be grouped into two categories. The first category encodes geometric and anatomical information. It consists of the UAC-projected three-dimensional vertex coordinates $(x,y,z)$ and the corresponding fibre direction vectors, yielding six spatial feature maps of size $(50,50)$. In addition, the total surface area of the left atrium is included as a global anatomical descriptor and broadcast across the spatial domain to form a constant $(1,50,50)$ grid. The second category encodes simulation-specific parameters. This includes the longitudinal and transverse conduction velocities $(\sigma_{l},\sigma_{t})$, which are broadcast across the grid to form a $(2,50,50)$ feature map, as well as the pacing location, represented as a spatial distribution on a $(1,50,50)$ grid. The operator is trained to predict the corresponding LAT map, represented on a $(1,50,50)$ grid.
Spatial Representation Learning.
To model spatial dependencies within the intrinsic coordinate domain, we employ a Vision Transformer (ViT) architecture [vit] to encode the multi-channel input grid into contextualised latent representations. Attention-based models have demonstrated strong capability for learning solution operators of partial differential equations by capturing long-range interactions and global context [li2023transformer, wu2024Transolver]. This capability is essential in our setting, where activation dynamics combine global wave propagation with sharp local transitions across the same spatial domain.
The input grid is first processed by a convolutional patch embedding layer [wu2021cvt], which partitions the domain into non-overlapping patches and projects each patch into a shared embedding space, yielding a sequence of latent tokens. Learnable positional embeddings are added to preserve spatial ordering within the tokenised representation. The token sequence is then processed by a stack of standard ViT blocks, each composed of multi-head self-attention and feed-forward sublayers, enabling the integration of global context and the modelling of long-range spatial interactions (Fig. 6c).
Following the encoder, the latent tokens are reshaped into a compact two-dimensional feature map. To recover full spatial resolution, we employ a lightweight decoder consisting of a transposed convolution layer for upsampling, followed by a convolutional projection head that maps latent features to the output field. The transposed convolution operation, originally introduced by [zeiler2011adaptive] for adaptive deconvolutional networks and later popularised for convolutional network visualisation by [zeiler2014visualizing], enables learnable upsampling through the transpose of the convolution operator. This decoding stage ensures that the global representations learned by the transformer are translated back into a high-resolution spatial field suitable for pixel-wise prediction.
Sobolev Loss with Spatial Smoothness Regularisations
@@
4.3 Instantiation for large deformation mechanics
We consider kinematic deformation of non-linear elastic solids under load as a representative volumetric mechanics problem. The computational domain is defined as a truncated ellipsoidal shell, a canonical approximation historically used to model left ventricular geometry [van1980application]. Although simplified, this geometry exhibits non-uniform curvature and provides a well-defined basal surface for the application of boundary conditions.
Inflation simulations were performed by fixing all degrees of freedom at the basal plane and applying a linearly increasing internal pressure to the inner surface. The resulting displacement field is governed by wall thickness, material stiffness, and applied pressure. The material behaviour was modelled using a hyperelastic transversely isotropic formulation based on the Guccione constitutive law [guccione1991passive]. Fibre directions were assigned circumferentially, with incompressibility enforced via a bulk modulus of $\kappa=1000,\mathrm{kPa}$. Fibre, cross-fibre, and fibre-sheet stiffness parameters were fixed at $b_{f}=b_{t}=b_{fs}=18.48$.
A dataset of 1,000 geometries was generated by varying four parameters: inner cavity diameter, wall thickness, bulk stiffness, and maximum internal pressure. Parameter ranges are reported in Table 1. All simulations were performed using the Cardiac Arrhythmia Research Package (CARP) [vigmond2003computational], with each simulation executed on 32 CPU cores.
For operator learning, displacement magnitude fields were projected from the physical mesh onto a shared intrinsic coordinate system defined by universal ventricular coordinates [bayer2018universal]. All fields were represented on a fixed $50\times 50\times 50$ grid in this coordinate domain and mapped back to the original geometry for evaluation.
Table 1: Inflation simulation parameter ranges.
Parameter Range
Inner cavity diameter (mm) 38–58 [lang2015recommendations]
Wall thickness (mm) 4–14 [kawel2012normal]
Bulk stiffness (dimensionless) 1–5
Maximum internal pressure (kPa) 1–10